# Detection and Characterization of Coordinated Online Behavior: A Survey

### v1

@@ 4.1.5. Content networks
A few works built and analyzed networks of content rather than users. After selecting the initial set of users, these methods build a fully connected network where the nodes are contents posted by the users (e.g., posts) and edge weights are proportional to the similarity between the linked contents. The community discovery process on a content network yields clusters of highly similar contents, regarded as proxies for coordination. User and content networks are interchangeable because each node in a user network can be mapped to the content they published, and each node in a content network can be mapped to its respective author, allowing for a dual representation of the same underlying actions. The main characteristics of the works based on content networks are reported in Table 6. Some works built and analyzed content networks where the nodes were Twitter cashtags [23, 24] or hashtags [26, 139]. Edge weights were given by the number of co-occurrences of the cashtags and hashtags in the same tweets. The analyses revealed suspicious clusters of contents that were later linked to online manipulations by coordinated actors. Lee et al. [67] created a network of similar texts published by multiple users, identifying clusters through connected components and maximal clique analysis. Similarly, [134] constructed two content networks: one for copypasta tweets, clustering coordinated actors via hierarchical clustering, and another to analyze the temporal evolution of co-occurring hashtags. Finally, [96] built an image similarity network by computing image embeddings and using Euclidean distance for comparison, then applied Louvain to detect groups of similar images.
@@ 4.2. Data mining and machine learning methods
Methods in this category follow the typical knowledge discovery in databases (KDD) analytical process, involving data collection and preparation, application of data mining and machine learning techniques, validation of the discovered patterns, and extraction of insights by interpretation of the results. The types of analyzed data include texts [5, 4, 104, 117, 108, 8], images [117], audio [84], interactions [113, 143, 119, 144], and temporal data [9, 62, 43]. In terms of machine learning techniques, some works leverage unsupervised approaches and focus on identifying groups of users exhibiting similar behaviors, while others rely on the availability of labeled data and apply supervised techniques to classify each user as either coordinated or non-coordinated. A comprehensive overview of these methods is presented in Table 7.
@@ 4.2.1. Unsupervised
Unsupervised methods work with unlabelled data and, therefore, do not exploit any knowledge on the membership of users to coordinated groups. Multiple unsupervised methods [4, 5, 104, 117] are based on the textClust [17] stream clustering algorithm. textClust is designed to group similar textual documents into micro-clusters that represent the topics recently discussed in the stream. The detection of rapidly growing clusters of documents is used to identify inorganic or orchestrated campaigns by coordinated users. This approach is conceptually similar to the analysis of a content network of similar documents that includes a time-based filter. A similar approach is also used in [8], where text and node embeddings are clustered via DBSCAN. Keller et al. [62] modeled user daily tweeting activity with a binary matrix whose cells $x_{jt}=1$ represent a user $u_{j}$ who tweeted at least once on day $t$, while cells $x_{jt}=0$ indicate no tweeting on that day. Then, groups of users with similar tweeting behaviors are found via expectation-maximization. Here, highly similar tweeting behaviors are considered as a proxy for coordination. Others modeled the sequence of user activities as temporal point processes. Sharma et al. [113] proposed the AMDN-HAGE generative model to jointly account for user activities and hidden group behaviours. The Attentive Mixture Density Network (AMDN) component models observed activity traces as a temporal point process, while the Hidden Account Group Estimation (HAGE) component models user groups as mixtures of multiple distributions. The synchronized groups of accounts detected by AMDN-HAGE are deemed coordinated, and the work makes the assumption that all detected coordinated groups are malicious. Similarly, Zhang et al. [143] jointly learned a distribution of user-group assignments based on how consistent each assignment is to the user embedding space and to some prior knowledge such as temporal logic. Finally, they used expectation-maximization to cluster the users according to the different distributions.
@@
Table 7. Data mining and machine learning methods for detecting coordinated behavior. For each group of works we report the input types, the machine learning approach, the learning paradigm, and whether the method takes time into account.
reference input machine learning approach learning time
[…]
[84] metadata, audio transcripts, thumbnails ensemble classification supervised
[…]
@@
In addition to the previous works, others proposed simpler methods or indicators as signals of possible coordinated behavior. Bellutta and Carley [9] analyzed account creation times, given that accounts involved in coordinated malicious activities are typically created in bursts [120]. They computed the daily histogram of account creations to which they applied a burst detection algorithm, identifying spikes in account creations by comparing the number of accounts created in a given day against the average number of daily accounts created in a reference time window. Another simple unsupervised technique is proposed in [119], where coordination is measured as the extent to which users converge—spontaneously or in an organized fashion—on the use of certain hashtags. The Gini coefficient, an indicator of inequality, is applied to the distribution of used hashtags to create time series where saddles close to 0—indicative of low inequality—correspond to no coordination whereas peaks close to 1—indicative of a situation where all users use the same few hashtags—correspond to strong coordination. Two alternative techniques are proposed in [81] and [82]. The former leverages contrast pattern mining to extract anomalous behavior, while the latter uses convergent cross mapping to discover cause-and-effect relationship and to construct an influence network. Similarly, [33, 34] use a discrete-time stochastic model to analyze coordinated activity, representing the user behaviors as interacting Markov chains.
@@ 4.2.2. Supervised
Supervised methods require labeled input data that they use to learn a function for predicting whether users belong to a coordinated group.  Zhang et al. [144] tackles a binary classification task to distinguish between coordinated and non-coordinated users that participate in cross-platform campaigns. They leverage information from an aid platform where coordinated users are known, to detect unknown coordinated users on a target platform. Input data consists of a known coordinated activity set on the aid platform, plus unknown user activities on the target platform. The relationship between the two activities are modeled with neural time series encoders before being fed to a multi-layer perceptron for prediction. Mariconti et al. [84] developed a system to identify YouTube videos that are likely to be raided (i.e., targeted by coordinated hate attacks). They used information coming from metadata, audio transcripts, and video thumbnails to compute machine learning features for each video. The prediction is given by an ensemble classifier that determines the likelihood that a new video will be raided based on a ground-truth of previously raided videos. A subsequent study focused on attributing coordinated hate attacks to the communities that organized them [108]. The system uses a peak detector to identify an abrupt rise in the comment activity of a YouTube video, a signal of a coordinated attack. Then, it leverages a trained classifier based on linguistic features from the comments to the video and a set of Reddit and 4chan communities, identifying the community responsible for each attack based on linguistic patterns and similarities. A simpler approach is proposed in [43] where a small ground-truth of 20 organic and inorganic campaigns is built by some subject-matter experts. Each campaign—the ground-truth ones and others to analyze—is characterized with indicators (i.e., features) belonging to three dimensions of participant behavior: network, time, and semantics. Then, [43] automatically labels as coordinated and inorganic those unknown campaigns whose indicators lay within the 95% confidence interval of ground-truth inorganic campaigns and outside the 95% confidence interval of ground-truth organic or non-coordinated campaigns.
@@ 4.3.1. Modeling complex coordination
defining dimensions reference concept indicators auth. harm. orch. time other user [5, 9, 18, 26, 51, 56, 90, 91, 92, 99, 101, 102, 139, 121, 27, 88] automation bot scores [56, 68, 67, 91, 99, 109, 121, 53, 141, 27, 113, 143, 88] moderation suspended users [91] username diversity entropy [109] activity account creation burstiness [14, 24, 26, 63, 67, 109, 110, 131, 135, 139, 45, 106, 133, 119, 62, 60] activity number of posts, retweets, … [106, 133] engagement number of views, likes, … [18, 101, 12, 60] timings action time interval content [14, 9, 26, 94, 91, 89, 133, 143, 135, 138, 139, 108, 24, 99, 62, 113, 143, 28, 88, 82] socio-linguistics attitudes, concerns, emotions, stances, topics, words [138, 139, 89, 60] repetitiveness text similarity scores [134, 122] socio-linguistics topic and hashtag evolution [15, 48] news diversity entropy, Gini coefficient [135, 47, 9] news reliability suspended and blacklisted URLs [106, 121, 8, 9] news reliability NewsGuard and MBFC scores [56] manipulation propaganda scores [70, 108] offensiveness toxicity scores [70, 99, 66, 121, 122] political bias MBFC scores, hashtag bias [18, 56, 70, 90, 91, 89, 99, 121, 19, 66, 88] coordination edge weights network [92, 48, 52, 66, 20, 29, 101, 64, 89, 27, 133, 88] influence centrality scores [24, 99, 19, 88] homophily assortativity [64, 96, 99, 48, 52, 92, 66, 89, 19, 27, 88] cohesiveness clustering coefficient, modularity, density [122, 68] stationarity user influx and outflux, user evolution $\dagger$: Media Bias/Fact Check
@@ 5. Characterization of coordinated behavior
When the coordination detection method is not integrated into an automated decision-making system, its output may initiate the characterization task. The aim of this task is to describe each coordinated user, group, or community along one or more of the defining dimensions outlined in Section 2.5. Characterizing the detected instances of coordination can also provide valuable information for validating detection results, especially in the absence of ground truth data. Characterizing coordinated online behavior is a semi-automatic task, as it involves some degree of manual analyses and observations by human analysts supported by some automatically-computed indicators that provide information on the coordinated actors. Table 8 summarizes the main indicators used for characterizing coordinated actors, whose discussion is presented in the remainder of this section. Indicators in table are grouped based on the information they leverage, which can be related to (i) users, (ii) content, or (iii) networks. The table also highlights the dimensions of coordination that each indicator addresses. Full dots indicate dimensions where the indicator has already been applied, while empty dots denote dimensions where it has potential use that has not yet been explored.
@@ 5.1. Authenticity
Authenticity refers to the extent to which users, groups, or communities correctly represent themselves to others on a platform. As a consequence, inauthenticity is found when an actor misrepresents itself, such as to mislead others on who they are or what they do. In other words, authenticity is a property of the coordinated actors. In literature, the most common proxy for inauthenticity is an account’s degree of automation, obtained via a bot score [5, 9, 18, 26, 51, 56, 90, 91, 92, 99, 101, 102, 139, 121, 27, 88]. By definition, social bots are accounts that make use of some or full automation [22]. Thus, accounts with a high bot score are likely to be mostly automated. Unfortunately however, while useful, the use of bot scores as indicators of inauthenticity faces some limitations. First, computing bot scores is a challenging task per se, and it is prone to errors [25]. Second, not all bots are inauthentic, as there exist some types of self-declared bots that operate for neutral or benign purposes [22]. Finally and most importantly, there exist multiple types of users that are inauthentic but that do not make use of automation (e.g., trolls, dissidents). Therefore, using bot scores as indicators of inauthenticity can cause both false positive and false negative errors. In addition to bot scores, some also used username similarity [91] and bursts of account creations [109] as proxies for inauthenticity. Table 8 also shows that the few existing indicators of inauthenticity are all based on user information. While this is expected since authenticity is a property of the actors rather than their actions, we also note that extremely short time intervals between user actions [18, 101, 12, 60] could be used as a red flag of automation and, by extension, also of inauthenticity.

### v2

@@ 4.1.5. Content networks
A few works built and analyzed networks of content rather than users. After selecting the initial set of users, these methods build a fully connected network where the nodes are contents posted by the users (e.g., posts) and edge weights are proportional to the similarity between the linked contents. The community discovery process on a content network yields clusters of highly similar contents, regarded as proxies for coordination. User and content networks are interchangeable because each node in a user network can be mapped to the content they published, and each node in a content network can be mapped to its respective author, allowing for a dual representation of the same underlying actions. The main characteristics of the works based on content networks are reported in Table 6. Some works built and analyzed content networks where the nodes were Twitter cashtags [23, 24], or hashtags [26, 146]. Edge weights were given by the number of co-occurrences of the cashtags and hashtags in the same tweets. The analyses revealed suspicious clusters of contents that were later linked to online manipulations by coordinated actors. Other works created a network of similar texts [64, 5] or comments [1] published by multiple users. Lee et al. [64] identify clusters through connected components and maximal clique analysis. Similarly, [141] constructed two content networks: one for copypasta tweets, clustering coordinated actors via hierarchical clustering, and another to analyze the temporal evolution of co-occurring hashtags. [99] built an image similarity network by computing image embeddings and using Euclidean distance for comparison, then applied Louvain to detect groups of similar images.
@@ 4.2. Data mining and machine learning methods
Methods in this category follow the typical knowledge discovery in databases (KDD) analytical process, involving data collection and preparation, application of data mining and machine learning techniques, validation of the discovered patterns, and extraction of insights by interpretation of the results. The types of analyzed data include texts [7, 6, 107, 123, 113, 11, 124], images [123], audio [83], interactions [118, 155, 126, 156, 136, 79, 80, 56, 154], and temporal data [12, 58, 40]. In terms of machine learning techniques, some works leverage unsupervised approaches (see Table 7) and focus on identifying groups of users exhibiting similar behaviors, while others rely on the availability of labeled data and apply supervised techniques (see Table 8). These supervised approaches can be further distinguished by their prediction target, namely user (individual account classification), community (groups of users), network (entire interaction graphs), target (entities targeted by coordinated actions).
@@ 4.2.1. Unsupervised
Unsupervised methods work with unlabelled data and, therefore, do not exploit any knowledge on the membership of users to coordinated groups. Multiple unsupervised methods [6, 7, 107, 123, 124] are based on stream clustering algorithms, designed to group similar textual documents into micro-clusters that represent the topics recently discussed in the stream. The detection of rapidly growing clusters of documents is used to identify inorganic or orchestrated campaigns by coordinated users. This approach is conceptually similar to the analysis of a content network of similar documents that includes a time-based filter. A similar approach is also used in [11], where text and node embeddings are clustered via DBSCAN. In [120], hashtags used by different users are leveraged to perform clustering via a Bayesian model. In contrast, [136, 63] apply clustering on interaction networks, such as the mention network [63], or more comprehensive networks incorporating mentions, replies, retweets, textual content, and hashtags [136]. Keller et al. [58] modeled user daily tweeting activity with a binary matrix whose cells $x_{jt}=1$ represent a user $u_{j}$ who tweeted at least once on day $t$, while cells $x_{jt}=0$ indicate no tweeting on that day. Then, groups of users with similar tweeting behaviors are found via expectation-maximization. Here, highly similar tweeting behaviors are considered as a proxy for coordination. Others modeled the sequence of user activities as temporal point processes. Sharma et al. [118] proposed the AMDN-HAGE generative model to jointly account for user activities and hidden group behaviours. The Attentive Mixture Density Network (AMDN) component models observed activity traces as a temporal point process, while the Hidden Account Group Estimation (HAGE) component models user groups as mixtures of multiple distributions. The synchronized groups detected by AMDN-HAGE are deemed coordinated and assumed to be malicious. Similarly, Zhang et al. [155] jointly learned a distribution of user-group assignments based on how consistent each assignment is to the user embedding space and to some prior knowledge such as temporal logic. Finally, they used expectation-maximization to cluster the users according to the different distributions.
@@
Table 7. Data mining and machine learning methods, based on unsupervised learning for detecting coordinated behavior. For each group of works we report the input types, the machine learning approach, and whether the method takes time into account.
reference input machine learning approach time
[6, 7, 107, 124] text streams text stream clustering
[123] text streams, image captions text stream clustering
[11] text, user-content network clustering of text and node embeddings
[58] daily tweeting activity expectation-maximization
[126] hashtags peak detection
[12] account creation timestamps burst detection
[79] user activities contrast pattern mining
[80] user activities convergent cross mapping
[33, 34] URL, hashtag, image, mention networked Markov chains
[118] user activities temporal point processes, gaussian mixture models
[155] user activities temporal point processes, expectation-maximization
[56] user activities Petri net
[154] user activities outlier detection
[55] text, posting time text similarity, timings of campaign launch and posting tweets
[120] hashtags bayesan model
[136] user activities cluster interactions networks
[63] mention cluster interactions networks
Table 8. Data mining and machine learning methods, based on supervised learning, for detecting coordinated behavior. For each group of works we report the input types, the machine learning approach, whether the method takes time into account, and the prediction target.
reference input machine learning approach time target
[156] user activities representation learning, conditional embedding, neural encoding user
[40] network, temporal, semantic features outlier detection network
[113] text peak detection, multiclass classification community
[108] user activities classifier user, target
[83] metadata, audio transcripts, thumbnails ensemble classification target
[4] user activities random weighted walk network
[84] co-retweet network graph neural network community
[57] retweet retrieval-augmented generation network
[88] network, text graph neural network user
@@
Other works proposed simpler methods or indicators as signals of possible coordinated behavior. Bellutta and Carley [12] analyze account creation times under the hypothesis that accounts involved in coordinated malicious activities tend to be created in bursts. They computed the daily histogram of account creations to which they applied a burst detection algorithm, identifying spikes in account creations by comparing the number of accounts created in a given day against the average number of daily accounts created in a reference time window. Another simple unsupervised technique is proposed in [126], where coordination is measured as the extent to which users converge—spontaneously or in an organized fashion—on the use of certain hashtags. The Gini coefficient, an indicator of inequality, is applied to the distribution of used hashtags to create time series where saddles close to 0—indicative of low inequality—correspond to no coordination whereas peaks close to 1—indicative of a situation where all users use the same few hashtags—correspond to strong coordination. Two alternative techniques are proposed in [79] and [80]. The former leverages contrast pattern mining to extract anomalous behavior, while the latter uses convergent cross mapping to discover cause-and-effect relationship and to construct an influence network. Similarly, [33, 34] use a discrete-time stochastic model to analyze coordinated activity, representing the user behaviors as interacting Markov chains. In [56] social media interactions are modeled as Petri nets, while in [154] coordination is detected by identifying anomalies in account sharing behavior. In [127] an exploratory analysis is conducted on YouTube links shared on 4chan.
@@ 4.2.2. Supervised
Supervised methods can be categorized based on their prediction target, leading to approaches that operate at the user, community, network, and target levels.
User.  Zhang et al. [156] tackles a binary classification task to distinguish between coordinated and non-coordinated users that participate in cross-platform campaigns. They leverage information from an aid platform where coordinated users are known, to detect unknown coordinated users on a target platform. Input data consists of a known coordinated activity set on the aid platform, plus unknown user activities on the target platform. The relationship between the two activities are modeled with neural time series encoders before being fed to a multi-layer perceptron for prediction. In [108] traditional classifiers (e.g., Random Forest) are employed on user- and content-level engagement features to address to detect both coordinated users and content targeted by coordination. In [88], a similarity network is built following [70], and multimodal node embeddings—combining one-hot features and text—are fed into a GNN to classify users as information operation drivers or legitimate accounts.
Community. Saeed et al. [113] focused on attributing coordinated hate attacks to the communities that organized them. A peak detector identifies an abrupt rise in the comment activity of a YouTube video, a signal of a coordinated attack. Then, it leverages a trained classifier based on linguistic features from the comments to the video and a set of Reddit and 4chan communities, identifying the community responsible for each attack based on linguistic patterns and similarities. In [84], a graph neural network model is exploited to recognize normal and abnormal communities.
Network. A simpler approach is proposed in [40], where a small expert-labeled ground truth of organic and inorganic campaigns is used to characterize campaigns along network, temporal, and semantic dimensions. Unknown campaigns are then labeled as coordinated if their indicators fall within the 95% confidence interval of inorganic campaigns and outside that of organic ones. In [4] a graph classification method based on random weighted walks leverages the density of local network structures to distinguish coordinated from non-coordinated networks. Kanakaris et al. [57] model retweet networks as propagation trees, encode them into text via graph prompting, and combine them with content and similarity-based examples in a RAG framework, enabling an LLM to classify coordinated disinformation campaigns.
Target. In [83], authors propose a system to predict YouTube videos likely to be targeted by coordinated hate attacks, using features from metadata, transcripts, and thumbnails, and an ensemble classifier trained on previously raided videos. As previously discussed, the approach in [108] operates at both the user and target levels.
@@ 4.3.1. Modeling complex coordination
4.4. Validation
The scarcity of labeled data limits the development of coordinated behavior detection methods. This constraint hampers the adoption of supervised approaches, contributing to the predominance of unsupervised techniques, and complicates both model tuning and evaluation. As a result, many methods are still assessed through post-hoc analyses rather than standardized quantitative benchmarks. To address this challenge, existing works adopt three main validation strategies.
Characterization-based validation. A common approach consists in validating methods through qualitative or descriptive analyses of the detected behaviors. In this setting, outputs are assessed a posteriori by examining account activity, temporal patterns, content and socio-linguistic signals, and network structure, to verify their consistency with known characteristics of coordinated behavior. While this strategy enables exploratory insights, it lacks objective ground truth and limits reproducibility and comparability across methods. Section 5 is dedicated to the task of characterization.
Public ground-truth datasets. Publicly available datasets can act as ground truth proxies, with some studies constructing balanced datasets from authoritative repositories [^6] https://transparency.x.com/en/reports/moderation-research.html (accessed: 31/07/2024) enabling both training [70] and evaluation [138, 91, 99, 18, 81, 52, 67, 104, 88, 108, 120, 56, 154, 79, 80, 118, 155] of detection methods. However, these datasets are often limited to specific platforms and coordination types, restricting model generalization.
Simulation-based validation. An alternative strategy consists in evaluating detection methods in controlled simulated environments, where coordinated behaviors can be explicitly modeled. Early approaches rely on agent-based models grounded in empirical observations. For instance, Mehta et al. [86] develop a multi-agent simulation of Reddit to analyze the impact of coordinated activity on recommendation systems, while Jahn et al. [53] design an agent-based model with heterogeneous agents (e.g., authentic users and coordinated boosters) to generate synthetic data. Other approaches generate synthetic datasets by injecting controlled anomalous patterns to enable systematic evaluation across coordination scenarios [158, 107]. Recent work explores large language models (LLMs) to simulate more realistic and adaptive behaviors. For instance, in [101] LLM-driven agents are used to mimic both organic users and coordinated actors under different regimes, from implicit alignment to explicit collaboration. Overall, simulation-based validation offers a flexible framework, though current efforts remain limited in scale and standardization.
@@ 4.4. Validation
defining dimensions
reference concept indicators auth. harm. orch. time other
user [7, 12, 17, 26, 47, 51, 93, 94, 95, 100, 102, 103, 146, 128, 28, 90, 92, 158] automation bot scores
[51, 65, 64, 94, 100, 115, 128, 49, 152, 28, 118, 155, 90] moderation suspended users
[94] username diversity entropy
[115, 13, 84, 61] activity account creation burstiness
[15, 24, 26, 59, 64, 115, 116, 138, 142, 146, 41, 110, 140, 126, 58, 55] activity number of posts, retweets, …
[110, 140, 136] engagement number of views, likes, …
[17, 102, 14, 55, 5, 150, 61] timings action time interval
content [15, 12, 26, 97, 94, 91, 140, 155, 142, 145, 146, 113, 24, 100, 58, 118, 155, 29, 90, 80, 13, 20, 122, 149, 136, 61] socio-linguistics attitudes, concerns, emotions, stances, topics, words
[145, 146, 91, 55] repetitiveness text similarity scores
[141, 129] socio-linguistics topic and hashtag evolution
[16, 44] news diversity entropy, Gini coefficient
[142, 43, 12, 20] news reliability suspended and blacklisted URLs
[110, 128, 11, 12, 20, 122, 150] news reliability NewsGuard and MBFC scores
[51, 84] manipulation propaganda scores
[66, 113] offensiveness toxicity scores
[66, 100, 62, 128, 129] political bias MBFC scores, hashtag bias
[20] fake content AI-generated
[17, 51, 66, 93, 94, 91, 100, 128, 18, 62, 90, 84] coordination edge weights
network [95, 44, 48, 62, 19, 30, 102, 60, 91, 28, 140, 90, 63] influence centrality scores
[24, 100, 18, 90, 149] homophily assortativity
[60, 99, 100, 44, 48, 95, 62, 91, 18, 28, 90, 122] cohesiveness clustering coefficient, modularity, density
[129, 65] stationarity user influx and outflux, user evolution
$\dagger$: Media Bias/Fact Check
@@ 5. Characterization of coordinated behavior
When the coordination detection method is not integrated into an automated decision-making system, its output may initiate the characterization task. The aim of this task is to describe each coordinated user, group, or community along one or more of the defining dimensions outlined in Section 2.5. Characterizing the detected instances of coordination can also provide valuable information for validating detection results, especially in the absence of ground truth data. Characterizing coordinated online behavior is a semi-automatic task, as it involves some degree of manual analyses and observations by human analysts supported by automatically computed indicators that provide information on the coordinated actors. Table 9 summarizes the main indicators used for characterizing coordinated actors, whose discussion is presented in the remainder of this section. Indicators in the table are grouped based on the information they leverage, which can be related to (i) users, (ii) content, or (iii) networks, a distinction that has also been adopted in related analyses of bots [96]. The table also highlights the dimensions of coordination that each indicator addresses. Full dots indicate dimensions where the indicator has already been applied, while empty dots denote dimensions where it has potential use that has not yet been explored.
@@ 5.1. Authenticity
Authenticity refers to the extent to which users, groups, or communities correctly represent themselves to others on a platform. As a consequence, inauthenticity is found when an actor misrepresents itself, such as to mislead others on who they are or what they do. In other words, authenticity is a property of the actors. In literature, the most common proxy for inauthenticity is an account’s degree of automation, obtained via a bot score [7, 12, 17, 26, 47, 51, 93, 94, 95, 100, 102, 103, 146, 128, 28, 90, 92, 158]. By definition, social bots are accounts that make use of some or full automation [25]. Thus, accounts with a high bot score are likely to be mostly automated. Unfortunately however, while useful, the use of bot scores as indicators of inauthenticity faces some limitations. First, computing bot scores is a challenging task per se, and it is prone to errors [96]. Second, not all bots are inauthentic, as there exist some types of self-declared bots that operate for neutral or benign purposes [25]. Finally and most importantly, there exist multiple types of users that are inauthentic but that do not make use of automation (e.g., trolls, dissidents). Therefore, using bot scores as indicators of inauthenticity can cause both false positive and false negative errors. In addition to bot scores, some also used username similarity [94] and bursts of account creations [115, 13, 84, 61] as proxies for inauthenticity. Table 9 also shows that the few existing indicators of inauthenticity are all based on user information. While this is expected since authenticity is a property of the actors rather than their actions, we also note that extremely short time intervals between user actions [17, 102, 14, 55, 5, 150, 61] could be used as a red flag of automation and, by extension, also of inauthenticity.